OpenStudy (anonymous):

Six times a number increased by seven more than the number is 42. Find the number

OpenStudy (anonymous):

So first we need to put this into a better form. (equation) Do you know how to do that?

OpenStudy (anonymous):

@Maddiem

OpenStudy (anonymous):

x * 6 + x + 7 = 42 Solve for x

OpenStudy (anonymous):

I think the equation is 6x+7x=42 But I don't think I'm correct 😳

OpenStudy (anonymous):

@BigPoppa117

OpenStudy (whpalmer4):

We'll let \(x\) be our number. "six times a number" -> \(6x\) "seven more than the number" -> \(7+x\) "is 42" -> \(= 42\) \[6x + 7 + x = 42\] can you solve that?

OpenStudy (anonymous):

Yes

OpenStudy (whpalmer4):

what do you get for the result?

OpenStudy (anonymous):

On sec

OpenStudy (anonymous):

You get x=6

OpenStudy (whpalmer4):

show me your work...

OpenStudy (whpalmer4):

or, try putting \(x=6\) into the equation, and tell me if it is true...

OpenStudy (whpalmer4):

\[6x+x+7 =42\]\[6(6)+6 + 7 = 42\]\[36 + 6 + 7 = 42\]\[36 + 13 = 42\]I don't think that is correct, do you?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

X=5

OpenStudy (anonymous):

Sorry I knew the answer was 5 I just screwed up thinking 35 divided by 7 was 6. My bad. I'm sorry

OpenStudy (whpalmer4):

first time I looked at the problem I concluded that \(x=7\) but then I always check my answers, and saved myself a bit of embarrassment :-)